OpenStudy (anonymous):

Which expression is equivalent to 3(5 + 6)? 8 + 9 8 + 6 15 + 18 15 + 6

OpenStudy (abdullahm):

The distributive property states: \(\sf \Huge a(b+c) = ab + ac\) So... \(\sf \Large 3 (5+6) = 3(5) + 3(6) = ?\)